OpenStudy (anonymous):

8x+2y=16 and -16x-7y=-20 using elimination find the x and y values

OpenStudy (leader):

8x+2y=16 -16x-7y=-20 Multiply the first equation by 2

OpenStudy (leader):

So we get 16x+4y=32 -16x-7y=-20 Now we just add both of the equations

OpenStudy (leader):

We get -3y=12 y=-4

OpenStudy (leader):

Now plug in y and solve for x.

OpenStudy (anonymous):

x=3

OpenStudy (leader):

8x+2(-4)=16 8x-8=16 8x=24 x=3

OpenStudy (leader):

So your answer will be (3,-4)

OpenStudy (leader):

Oh lol @username255 didn't realize you asked the question..

OpenStudy (anonymous):

um yeah it's ok thank you though

OpenStudy (leader):

Yeah If I helped you you can medal me and if you want you can fan me so that way if you need my help you can find my username in your fan list..

OpenStudy (leader):

This is how you give medals http://prntscr.com/55hgk5

OpenStudy (leader):

And this is hw you fan http://prntscr.com/55hgww For some reason it gives me mod powers... but they don't work

OpenStudy (leader):

@username255